If the earth is 1/4th of its present distance from the sun, the duration of the year would be
(a) 1/4 of the present year
(b) 1/6 of the present year
(c) 1/8 of the present year
(d) 1/16 of the present year

Grade:12

1 Answers

asjsaafsohfaohfsi
7 years ago
answer=c
t^2=(k)r^3
k is a constant
now
r’=r/4
but the value in the equaton
t’^2=kr’^2
solving we get
t’=t/8